OpenStudy (anonymous):

does x^2+y^2+z^2 equal 1?

OpenStudy (mayaal):

Yes I think so....!

OpenStudy (anonymous):

ohh okay perfect so if I have a a square root of (4x^2 + 4y^2 + 4z^2), then i can factor out the 4, and have x^2+y^2+z^2 =1, and then i will be left with sq.root(4)?

OpenStudy (ybarrap):

take, x=1, y=1, z=1, then x^2+y^2+z^2=3

OpenStudy (anonymous):

no but there aren't any given values for x y and z

OpenStudy (ybarrap):

So, generally this equation is not true, except for some sets of x,y and z. So generally the equation does not equal 1

OpenStudy (anonymous):

ohhhhh okay. i was thinking of how generally, sin^2+cos^2=1 so i thought this might work for my equation...

OpenStudy (ybarrap):

yes, it reminds us of that equation, but that equation is justified because it is the hypotenuse of the unit circle is always 1, x^2 + y^2 + z^2 = 1 is a sphere of unit radius. Not all spheres have unit radii.

OpenStudy (anonymous):

oohhh okay thanks then! :-)

OpenStudy (ybarrap):

good luck

OpenStudy (anonymous):

thanks I need all the luck i can get. lol have a calc exam next week :-(